# A Survey on Integrated Sensing, Communication, and Computation

### v1

@@ VI-B3 Coexistence of ISC and MEC Tasks
VI-C Task-oriented ISCC Resource Allocation
@@
In this paradigm, network resources are allocated among the sensing, communication, and computation modules of a complex task to achieve its customized goal. The fulfillment of complex tasks requires the cooperation of sensing for information acquisition, communication for information transmission, and computation for information processing, such as auto-driving tasks, federated learning tasks, edge inference tasks, etc. The qualities of the three modules decide the performance of these tasks and they compete for the limited radio resources. Moreover, there is a lack of a uniform design objective for these tasks. For example, instead of targeting throughput maximization in conventional communication tasks, a federated learning task concerns improving learning accuracy while reducing training latency. Therefore, various kinds of task-oriented ISCC resource allocation schemes are developed, as discussed below.
VI-C1 ISCC for Vehicular Networks
@@
In this case, task-oriented ISCC resource allocation schemes are proposed to enhance the system performance, e.g., achieve low latency for auto-driving [89] and address road traffic congestion [294]. Specifically, in [89], the message and resource allocation are investigated for cooperative perception in fog-based vehicular networks. The satisfaction of cooperative perception (SCP) is firstly characterized based on the spatial-temporal value and latency performance. To maximize the sum SCP, the sensing block message, communication resource block, and computation resource are jointly allocated under the maximum latency and sojourn time constraints of vehicles. Besides, in [294], the ISCC technique is regarded as an appealing control tool for road traffic congestion. The evolution of road traffic congestion control is divided into two sequential stages, i.e., the sensing and the communication stages. In the former stage, congestion control strategies experience rapid growth due to the accessibility of traffic data. In the latter stage, with advances in vehicular networks, big data and AI, the improvement of the ability to collect finer data is significant. This paper also introduces how existing congestion detection techniques evolve with the continuous development of sensing, communication, and computation capability in the networks.
VI-C2 ISCC for Machine Communications
In on-purpose machine communication (OMC), the interactive information flow is the machine semantics, which refers to the consensus information reached by the intelligent machines (IMs) after feature extraction and transformation from the raw data [90]. A cross-layer ISCC framework is designed for OMC in [295], where the ISCC is deployed at the cloud, edge, and terminals, and the IMs can transmit directional beams to the dedicated devices according to their communication purposes
VI-C3 ISCC for Mobile Crowdsensing
@@
In [75], mobile crowdsensing is modeled as a typical sensing, computation, and communication process. Due to hostile wireless channels, limited spectrum, energy, and computation resources, the transmission or processing of sensing data may be unsuccessful in wireless edge networks. Thus, an ISCC framework for multi-dimensional resource-constrained crowdsensing systems is proposed to optimize the overall performance by jointly selecting users, allocating bandwidth, and designing the strategies.
VI-C4 ISCC for Sensor Networks
@@
VI-C5 ISCC for Federated Learning Tasks
In [48], the online federated-learning-guided ISCC schemes are proposed. The sensing, on-device computation, and communication operate in a time-division manner. The FMCW radar signals are transmitted by the MDs in the sensing mode. The radar echo signals are received by the MDs and locally processed to extract the target information in the computation mode. The locally processed data is then delivered from MDs to the server for global model updating in the communication mode. Aim at maximizing the convergence speed of model training, the sensing, communication, and computation modules are carefully designed. However, the competition for time resources will limit the performances of these modules and thus the learning performance.
VI-C6 ISCC for Edge AI Inference Tasks
For real-time edge inference tasks, a task-oriented AirComp scheme is first proposed in [26]. In this work, the AI model is split into two parts, the front-end part deployed at devices is used for feature extraction from sensory data. The backend part with intensive computation deployed at an edge server completes the inference task by receiving all local feature vectors from all devices. Specifically, multiple devices sense the same wide view of the source target and obtain a local distortion-corrupted version of the ground-true sensory data. Then, each device extracts a low-dimensional local feature vector from its local sensory data using a principal component analysis based linear extractor. An edge server aggregates all local feature vectors via AirComp to obtain a denoised global feature vector, which is input into the remaining AI model for finishing the downstream inference task. This work directly adopts the inference accuracy as the design objective rather than the conventional metric of MMSE, as the latter ignores the feature elements with the same distortion but heterogeneous contributions to the inference accuracy. Besides, the communication efficiency is enhanced using AirComp. The above work is further extended to the case with sensing noise suppression [296] and the decentralized case [297]. Besides, to deal with the scenario where different devices sense disjoint narrow views of the source target, a task-oriented sensing, feature quantization, and feature transmission technique is proposed in [47], where the local feature vectors of different devices are cascaded at the server to form a global one. Furthermore, a device scheduling based ISCC approach is proposed in [298] for enhancing the inference accuracy while preserving the data privacy.
VII ISCC in Future Advanced Networks
@@
In this section, the potential ISCC techniques for deploying digital twin platforms at the network edge and the potential ISCC applications in two future advanced networks, i.e., the computing power networks and the SAGINs, are discussed.
Figure 8: An ISCC-based digital twin network.
VII-A ISCC for Digital Twin Enabled Wireless Networks
@@
A digital twin system refers to constructing virtual or digital representations of the physical objects and systems via the techniques of high-quality distributed sensing, precise 3D map construction, real-time tracing, deep machine learning, and so on [299]. It targets generating simulation data that cannot be distinguished from its real-world counterparts, conducts data analysis, and makes decisions based on the generated data [300]. Digital twins have been widely used in many areas e.g., intelligent manufacturing, climate change, healthcare, and economic crises [300]. Particularly, since digital twins have a powerful ability to simulate complex wireless environments, it has attracted more and more attention to deploy digital twins at the network edge for addressing many issues including network complexity control [301], network performance enhancement [302], disruptions prediction and prevention [303], network topology visualization [304], and supporting MEC and industrial IoT networks [299].
@@
However, establishing a digital twin network is challenging. It desires massive sensory data of multiple modalities for acquiring ground-true information of real wireless networks, high communication throughput for transmitting sensory data or the corresponding extracted information, and powerful computation capability for data generation, analytics, and decision-making. To tackle these challenges, an ISCC-based digital twin network is proposed as shown in Fig. 8, where many sensors of different types are deployed for perceiving a real wireless network to obtain ground-true information, and the digital twin system is deployed at an edge server for executing all complex and computation-intensive algorithms to construct the virtual representation. Based on such a digital twin network, two ISCC techniques are discussed. One is a hybrid learning platform for training AI models that are used for constructing digital twin systems. The other is a real-time ISCC technique to leverage the established digital twin for making accurate and fast decisions.
VII-A1 Hybrid AI Training Based Digital Twin Network Construction
One significant step for establishing the digital twin network in Fig.8 is training many AI models for accomplishing many intelligent tasks, by using a large amount of multi-modal sensory data from many distributed sensors. However, different sensors have heterogeneous resources like energy and different capabilities in terms of sensing, communication, and computation. Besides, different sensors have different data privacy preservation requirements. To deal with this heterogeneous digital twin network, a hybrid AI training platform that integrates federated learning, federated split learning, and centralized learning, is proposed to meet various requirements of different sensors, as shown in Fig. 9. For example, for the sensors with critical privacy concerns and high computation and communication capabilities, federated learning is adopted. For the sensors with high privacy preservation requirements and low computation and communication abilities, federated split learning is performed. For the sensors without privacy concerns and with limited communication and computation resources, they can directly send the raw sensory data to the edge server for centralized learning.
@@ VII-A1 Hybrid AI Training Based Digital Twin Network Construction
Figure 9: An ISCC-based hybrid AI model training platform for establishing digital twin.
@@
VII-A2 ISCC Based Digital Twin Network Application
@@
VII-B Computing Power Networks Supported ISCC
@@
VII-B1 Computing Sharing for Various ISCC Tasks
@@
As shown in Fig. 10, an edge-device cooperative framework is deployed for supporting several ISCC tasks, where sensing and some basic sensory data processing for feature extraction are conducted at device and the deep neural networks based complicated computation tasks are performed at edge server. As a result, this framework can enjoy three-fold benefits of privacy preservation via avoiding raw sensory data transmission, low communication due to low-dimensional feature extraction, and computation offloading. In Fig. 10, two computing architectures are compared for completing multiple ISCC tasks. In the conventional monolithic computing architecture as shown in Fig. 10(a), each task has its own individual sensory data processing streams. The intermediate computing results cannot be shared among different tasks. Consequently, it leads to a severe waste of computing resources as well as increasing storage and communication costs. To tackle these issues, the micro-service architecture divides the signal processing of each task into many fine-granularity computing blocks as shown in Fig. 10(b). Different tasks can share the same computing blocks, e.g., the blocks of signal sampling, clutter cancellation, and feature extraction for wireless sensing signal and the blocks of stream processing, frame analysis and content identification for video. This significantly saves the computing and storage overhead as well as reduces the communication load for transmitting the on-device processing results to the edge server.
Figure 11: A directed graph model of computing power networks supported ISCC tasks.
@@
VII-B2 Directed-graph-based Joint Scheduling of Sensing, Communication, and Computation
@@
In computing power networks, the fine-granularity computing blocks for completing to make full utilization of the computing resources in various devices. Thereby, the completion of ISCC tasks requires cooperation of multiple nodes. For enhancing the performance of ISCC tasks while saving the resources in terms of sensing, computation, and communication, a directed graph model can be constructed as shown in Fig. 11. In the graph, there are two kinds of devices or called nodes. One is the sensing nodes with the capability of sensing, on-device computation, and communication. The other is the computing nodes with the ability of communication and computation. A target can be sensed by multiple sensing nodes and same computing blocks are deployed at different computing devices. As a result, the fulfillment of an ISCC task has multiple choices of paths. To be specific, a directed edge from node $a$ to node $b$ means that node $a$ transmits its computing results to node $b$ for downstream processing. In each path, sensing and all required computing blocks are sequentially performed for completing the task. Whether there is a directed edge between two nodes depends on whether a wireless communication link can be constructed between them for reliably transmitting the intermediate computing results based on the factors like transmit power, bandwidth, time, and channel gain. In order to enhance the task performance while saving the sensing, communication, and computation costs, it’s desirable to find the best path in the directed graph.
VII-C Space-Air-Ground Integrated Networks Supported ISCC

### v2

@@ VII-B3 Coexistence of ISC and MEC Tasks
As well as the above three cases, authors in [307] propose a more complex framework where an edge server simultaneously provides the ISCC-supported DNN splitting inference services and the communication services for multiple UAVs. The average sum rate of all UAVs is maximized under the latency and accuracy guarantees of the DNN splitting inference tasks.
VII-C Task-oriented ISCC Resource Allocation
@@
A massive number of intelligent devices are deployed at the network edge to complete various kinds of tasks for enabling intelligent connection of everything [308]. In this paradigm, network resources are allocated among the sensing, communication, and computation modules of a complex task to achieve its customized goal. The fulfillment of complex tasks requires sensing for information acquisition, communication for information transmission, and computation for information processing, such as auto-driving tasks, federated learning tasks, and edge inference tasks [12, 40]. In other words, it is desirable to generate and deliver the most valuable information that can be used to effectively and efficiently complete an intelligent computation task [308]. However, the qualities of the three modules decide the performance of these tasks but the three modules compete for the limited radio resources. Moreover, there is a lack of a uniform design objective for these tasks. For example, instead of targeting at throughput maximization in conventional communication tasks, a federated learning task concerns improving learning accuracy while reducing training latency. To this end, various task-oriented ISCC resource allocation schemes are developed, as discussed below.
VII-C1 ISCC for Federated Learning Tasks
In each training iteration of an online federated learning task, an edge server first broadcasts a global model to all participating devices. Then each device acquires training data samples via sensing, calculates the local gradient via computation, and uploads the local gradient to the server via communication [291, 275, 309]. The server finally aggregates all local gradients to update the global model. In this procedure, on-device sensing, computation, and communication are tightly coupled for the following reasons. The learning performance, e.g., convergence rate and per-iteration loss function degradation, depends on the per-iteration gradient distortion, which is decided by the number of obtained and processed data samples, the sensory data distortion (sensing SNR), and the transmission distortion (communication SNR). However, it is hard to enhance all these metrics due to the limited network resources [275]. To this end, the learning performance guided ISCC resource allocation schemes are desirable. In [291], by operating sensing, on-device computation, and communication in a time-division manner and using FMCW for sensing, an ISCC-based federated learning framework is proposed with the goal of accelerating the convergence rate. However, the influence of sensing noise on the convergence performance is ignored in [291], which limits the application of its proposed scheme especially in the case of hostile sensing channels and limited network resources. This scheme is further extended to the case of UAV-assisted federated learning [310]. To characterize the influence of sensing noise, the authors in [275] theoretically characterize the influence of sensory data distortion, communication SNR, and the number of sensed and processed data samples on the convergence rate. Guided by this analysis, an ISCC resource allocation scheme is proposed for over-the-air federated learning.
VII-C2 ISCC for Edge AI Inference Tasks
A real-time inference task involves sensing for data acquisition, communication for data sharing, and computation for data processing and intelligent decision-making [12]. The inference accuracy depends on the distortion caused during the three processes, e.g., sensing clutter and noise, quantization distortion, and channel noise, but the three processes compete for network resources to enhance their own qualities. Therefore, to enhance the inference accuracy with critical latency requirement, several ISCC resource management schemes are proposed. In [27], the AI model is split into two parts, the front-end part deployed at devices is used for feature extraction from sensory data. The backend part with intensive computation deployed at an edge server completes the inference task by receiving all local feature vectors from all devices. Specifically, multiple devices sense the same wide view of the source target and obtain a local distortion-corrupted version of the ground-true sensory data. Then, each device extracts a low-dimensional local feature vector from its local sensory data using a principal component analysis based linear extractor. An edge server aggregates all local feature vectors via AirComp to obtain a denoised global feature vector, which is input into the remaining AI model for finishing the downstream inference task. This work directly adopts the inference accuracy as the design objective rather than the conventional metric of MMSE, as the latter ignores the feature elements with the same distortion but heterogeneous contributions to the inference accuracy. Since the instantaneous inference accuracy lacks a mathematical model, a surrogate metric called discriminant gain is applied for classification tasks [311]. It measures the distances between different classes in the feature space based on the symmetric KL divergence. Particularly, with a higher value of discriminant gain, different classes are better separated in the feature space, leading to a higher achievable inference accuracy. Besides, the communication efficiency is enhanced using AirComp. The above work is extended to the case with sensing noise suppression [42], the decentralized case [312], and the case of unbalanced classification accuracy [313]. To further deal with the scenario where different devices sense disjoint narrow views of the source target, a task-oriented sensing, feature quantization, and feature transmission technique is proposed in [290], where the local feature vectors of different devices are cascaded at the server to form a global one. Based on [290], an ISCC-based selection approach is proposed in [314] to choose the best from the modes of on-device, on-server, and edge-device cooperative. Moreover, ISCC-based joint device scheduling and resource allocation methods are proposed for enhancing the inference accuracy while preserving the data privacy [315] and simultaneously completing multiple inference tasks [316].
VII-C3 ISCC for Vehicular Networks
@@
In this case, task-oriented ISCC resource allocation schemes are proposed to enhance the system performance, e.g., achieve low latency for auto-driving [317] and address road traffic congestion [318]. Specifically, in [317], the message and resource allocation are investigated for cooperative perception in fog-based vehicular networks. The satisfaction of cooperative perception (SCP) is firstly characterized based on the spatial-temporal value and latency performance. To maximize the sum SCP, the sensing block message, communication resource block, and computation resource are jointly allocated under the maximum latency and sojourn time constraints of vehicles. Besides, in [318], the ISCC technique is regarded as an appealing control tool for road traffic congestion. The evolution of road traffic congestion control is divided into two sequential stages, i.e., the sensing and the communication stages. In the former stage, congestion control strategies experience rapid growth due to the accessibility of traffic data. In the latter stage, with advances in vehicular networks, big data and AI, the improvement of the ability to collect finer data is significant. This paper also introduces how existing congestion detection techniques evolve with the continuous development of sensing, communication, and computation capability in the networks. Moreover, the sensing, communication, and computation are modeled in the scenario of continuous and random task arrival in vehicular networks and a collaborative sensory data fusion framework is proposed in [319] for maximizing the task completion rate.
VII-C4 ISCC for Mobile Crowdsensing
@@
In the paradigm of mobile crowdsensing, devices sense the environment to collect data via the equipped sensors and offload the data to a server after pre-processing to reduce the communication load for further processing. The design goal is to maximize a data utility function, which depends on the acquired data quantity and quality [169]. Due to hostile wireless channels, limited spectrum, energy, and computation resources, the transmission or processing of sensing data may be unsuccessful in wireless edge networks. To this end, an ISCC framework for multi-dimensional resource-constrained crowdsensing systems is proposed in [169] to optimize the overall performance by jointly selecting users, allocating bandwidth, and designing the strategies. It should be emphasized that the data utility function in this paradigm cannot well characterize the customized goals of the federated learning and edge AI inference tasks, although the sensing, computation, and communication procedures are similar. For example, the data utility function fails to quantify the distortions caused during sensing, computation, and communication on the learning or inference performance as well as the long-term influence on the convergence rate.
VII-C5 ISCC for Sensor Networks
@@
VII-C6 ISCC for Robotics
An ISCC-based unmanned UAV-robot system is proposed in [321], where the UAVs enabled with the functions of sensing, communication, and computation, are deployed for controlling robotics for task execution. Specifically, UAVs process the collected real-time sensory data to generate control commands, which are further transmitted to robotics for execution. Such workflows require the cooperation and coordination of sensing, computation, communication, and control (SC3). To this end, a task-oriented closed-loop SC3 optimization approach is proposed to simultaneously overcome the scarce channel bandwidth, limited energy supply, and sensing distortion for minimizing total linear quadratic regulator cost.
VII-D Discussion
By enabling resource sharing, coordination, and joint optimization among sensing, communication, and computation under multiple objectives or a customized goal, network resource utilization and task performance can be enhanced. However, the problems of RRM in ISCC systems are more complex than those of the separated designs, as more tightly coupled variables should be jointly optimized. Moreover, a customized scheme should be designed for different scenarios and tasks. The current ISCC RRM strategies have relatively low generalization capability, leading to high implementation costs in different settings. To this end, more general, simple and efficient ISCC RRM strategies are demanded in the future. On the other hand, ISCC RRM is suitable for critical tasks and cases with extreme performance requirements or limited network resources. Otherwise, the separated RRM designs are sufficient.
As well as the tasks mentioned previously, there are many other tasks that involve the coordination of sensing, communication, and computation to enhance performance with the limited network resources, e.g., holographic communication, auto-driving, intelligent agriculture, and smart healthcare. In these tasks, the sensing, communication, and computation processes are tightly coupled via competing network resources but cooperate to achieve a customized goal. However, the design of task-oriented ISCC RRM for these tasks demands further exploration.
VIII ISCC in the Future
@@
In this section, three future implementation examples of ISCC are first provided, including the potential ISCC techniques for deploying digital-twins platforms at the network edge and the potential ISCC applications in two future advanced networks, i.e., the computing power networks and the SAGINs. Then, several critical unresolved issues that are significant to efficiently implement ISCC in the future are discussed.
VIII-A ISCC for Digital-Twins Enabled Wireless Networks
Figure 11: An ISCC-based digital-twins network.
@@ VIII-A ISCC for Digital-Twins Enabled Wireless Networks
A digital-twins system refers to constructing virtual or digital representations of the physical objects and systems via the techniques of high-quality distributed sensing, precise 3D map construction, real-time tracing, deep machine learning, and so on [322]. It targets generating simulation data that cannot be distinguished from its real-world counterparts, conducts data analysis, and makes decisions based on the generated data [323]. Digital twins have been widely used in many areas e.g., intelligent manufacturing, climate change, healthcare, and economic crises [323]. Particularly, since digital twins have a powerful ability to simulate complex wireless environments, it has attracted more and more attention to deploy digital twins at the network edge for addressing many issues including network complexity control [324], network performance enhancement [325], disruptions prediction and prevention [326], network topology visualization [327], and supporting MEC and industrial IoT networks [322].
@@
However, establishing a digital-twins network is challenging. It desires massive sensory data of multiple modalities for acquiring ground-true information of real wireless networks, high communication throughput for transmitting sensory data or the corresponding extracted information, and powerful computation capability for data generation, analytics, and decision-making. To tackle these challenges, an ISCC-based digital-twin network is proposed as shown in Fig. 11, where many sensors of different types are deployed for perceiving a real wireless network to obtain ground-true information, and the digital-twins system is deployed at an edge server for executing all complex and computation-intensive algorithms to construct the virtual representation. Based on such a digital-twins network, two ISCC techniques are discussed. One is a hybrid learning platform for training AI models that are used for constructing digital-twins systems. The other is a real-time ISCC technique to leverage the established digital twins for making accurate and fast decisions.
@@
Figure 12: An ISCC-based hybrid AI model training platform for establishing digital twins.
VIII-A1 Hybrid AI Training Based Digital-Twins Network Construction
One significant step for establishing the digital-twins network in Fig.11 is training many AI models for accomplishing many intelligent tasks, by using a large amount of multi-modal sensory data from many distributed sensors. However, different sensors have heterogeneous resources like energy and different capabilities in terms of sensing, communication, and computation. Besides, different sensors have different data privacy preservation requirements. To deal with this heterogeneous digital-twins network, a hybrid AI training platform that integrates federated learning, federated split learning, and centralized learning, is proposed to meet various requirements of different sensors, as shown in Fig. 12. For example, for the sensors with critical privacy concerns and high computation and communication capabilities, federated learning is adopted. For the sensors with high privacy preservation requirements and low computation and communication abilities, federated split learning is performed. For the sensors without privacy concerns and with limited communication and computation resources, they can directly send the raw sensory data to the edge server for centralized learning.
@@ VIII-A1 Hybrid AI Training Based Digital-Twins Network Construction
VIII-A2 ISCC Based Digital-Twins Network Application
@@
VIII-B Computing Power Networks Supported ISCC
@@
VIII-B1 Computing Sharing for Various ISCC Tasks
@@
As shown in Fig. 13, an edge-device cooperative framework is deployed for supporting several ISCC tasks, where sensing and some basic sensory data processing for feature extraction are conducted at device and the DNNs based complicated computation tasks are performed at edge server. As a result, this framework can enjoy three-fold benefits of privacy preservation via avoiding raw sensory data transmission, low communication due to low-dimensional feature extraction, and computation offloading. In Fig. 13, two computing architectures are compared for completing multiple ISCC tasks. In the conventional monolithic computing architecture as shown in Fig. 13(a), each task has its own individual sensory data processing streams. The intermediate computing results cannot be shared among different tasks. Consequently, it leads to a severe waste of computing resources as well as increasing storage and communication costs. To tackle these issues, the micro-service architecture divides the signal processing of each task into many fine-granularity computing blocks as shown in Fig. 13(b). Different tasks can share the same computing blocks, e.g., the blocks of signal sampling, clutter cancellation, and feature extraction for wireless sensing signal and the blocks of stream processing, frame analysis and content identification for video. This significantly saves the computing and storage overhead as well as reduces the communication load for transmitting the on-device processing results to the edge server.
@@
VIII-B2 Directed-graph-based Joint Scheduling of Sensing, Communication, and Computation
@@
In computing power networks, the fine-granularity computing blocks for completing to make full utilization of the computing resources in various devices. Thereby, the completion of ISCC tasks requires cooperation of multiple nodes. For enhancing the performance of ISCC tasks while saving the resources in terms of sensing, computation, and communication, a directed graph model can be constructed. In the graph, there are two kinds of devices or called nodes. One is the sensing nodes with the capability of sensing, on-device computation, and communication. The other is the computing nodes with the ability of communication and computation. A target can be sensed by multiple sensing nodes and same computing blocks are deployed at different computing devices. As a result, the fulfillment of an ISCC task has multiple choices of paths. To be specific, a directed edge from node $a$ to node $b$ means that node $a$ transmits its computing results to node $b$ for downstream processing. In each path, sensing and all required computing blocks are sequentially performed for completing the task. Whether there is a directed edge between two nodes depends on whether a wireless communication link can be constructed between them for reliably transmitting the intermediate computing results based on the factors like transmit power, bandwidth, time, and channel gain. In order to enhance the task performance while saving the sensing, communication, and computation costs, it’s desirable to find the best path in the directed graph.
VIII-C Space-Air-Ground Integrated Networks Supported ISCC